OpenStudy (anonymous):

what is the apparent weight of 60kg man in a car accelerating at 5m/s^2 down a straight road. my equation W(app)=MA-MG or W(app)=M(a-g) doesnt seem to work

OpenStudy (anonymous):

its horizontal sorry

OpenStudy (anonymous):

So the acceleration is perpendicular to gravity.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

As acceleration due to gravity is 9.8ms^-2, the magnitude of the net acceleration, according to Pythagoras, is\[\sqrt{9.8^2+5^2}ms^{-2}\]Therefore apparent weight is\[60 \sqrt{9.8^2+25}N\]And it acts in a direction\[\arctan{\frac{5}{9.8}}\]from the vertical.

OpenStudy (anonymous):

oh ok so the it would look like this 

OpenStudy (anonymous):

Yup, that pretty much sums it up. Vector summation.

OpenStudy (anonymous):

ok that makes so much more sense thank you